OpenStudy (anonymous):

Prove: If n ≡ 1 (mod 6), then n2 + 2n is composite, where n>1. @MIT 18.02 Multiva…

OpenStudy (anonymous):

if \[n\equiv 1( \text{ mod }6)\] then \[n+2\equiv 3(\text{ mod } 6)\] and therefore \[n(n+2)\equiv 3(\text { mod } 6)\] meaning it is divisible by 3

OpenStudy (anonymous):

a more prosaic way of thinking about this is if \[n\equiv 1 ( \text { mod } 6)\] then \[n=1+6k\] so \[n+2=3+6k\] making \[n(n+2) = (1+6k)(3+6k)=3+24k+36k^2\] which is clearly divisible by 3

myininaya (myininaya):

\[n^2+2n=n(n+2)\] since this is factorable and is an integer isn't n(n+1) composite for all integers >1

myininaya (myininaya):

oops n(n+2)

OpenStudy (anonymous):

good point. now that i think about it the problem is rather silly

myininaya (myininaya):

you don't need the first part to prove it its already done

OpenStudy (anonymous):

just kidding

myininaya (myininaya):

first type i was just a answer checker for first problem so there

OpenStudy (anonymous):

first type of typo?

myininaya (myininaya):

lol

myininaya (myininaya):

of or or?

OpenStudy (anonymous):

omg. I just realized part of the question was typed wrong! Sry! it is supposed to be n^2 + 2^n :(

myininaya (myininaya):

i would use induction

myininaya (myininaya):

\[\text{ so we have } n=6k+1\] For k=1 we have n=8 \[8^2+2^8=64+256=320 \text{ this number is composite}\] now assume it is true for some integer k=j we hae n=6j+1 this means this is true \[(6j+1)^2+2^{6j+1}=(w)(r) \text{ for some integers w and r}\] now we want to show it is true for k=j+1 we have n=6(j+1)+1 so \[(6(j+1)+1)^2+2^{6(j+1)+1}=(6j+1+6)^2+2^{6j+1+6}\] \[(6j+1)^2+2(6j+1)(6)+(6^2)+2^{6j+1} \cdot 2^6\] \[(6j+1)^2+2^{6j+1} \cdot 2^6 +12(6j+1)+36\] .... thinking...

myininaya (myininaya):

i give up i can't do math anymore i can't even evaluate n for k=1

myininaya (myininaya):

which i know is n=7

OpenStudy (anonymous):

i think we might have to use that if p is prime then it is congruent to 1 or 5 mod 6

myininaya (myininaya):

i have to go do some stuff

OpenStudy (anonymous):

we know n^2 is congruent to 1 mod 6 because n is

OpenStudy (anonymous):

so it comes down to what is 2^n congruent to mod 6

OpenStudy (anonymous):

k , thanks myininaya

OpenStudy (anonymous):

the choices are only 2 and 4 because it is a power of 2

OpenStudy (anonymous):

so maybe we can show that it is not 4, in which case we are done

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ah yes and it is not 4, it is 2

OpenStudy (anonymous):

reason being that \[2^{1+6k}=2\times 2^{6k}\] and \[22^6=64\equiv 4(\text { mod } 6)\] and so (pretty sure of this but you should check \[2^{6k}\equiv 4 (\text { mod } 6)\] so \[2^{6k+1}\equiv 2( \text { mod } 6)\]

OpenStudy (anonymous):

sorry i meant \[2^6=64\equiv 4(\text{ mod } 6)\]

OpenStudy (anonymous):

ok. So then the sum of n^2 + 2^n , which I think would be n^2 + 2^(6k+1), is composite?

OpenStudy (anonymous):

i'm sry. I'm a little confused. :/

OpenStudy (anonymous):

yes it is composite because a prime number must be congruent to 1 or 5 mod 6

OpenStudy (anonymous):

n^2 is congruent to 1 mod 6 because n is

OpenStudy (anonymous):

and \[2^{1+6k}\] is congruent to 2 mod 6 so their sum is congruent to 3 mod 6 unless i made some mistake